Assume PI sta = 18+00, PI elev. = 100.00, L = 18 stations, incoming grade = +2%, outgoing grade = -4%
1.  The elevation at station 25+00 is
(a) 71.33
(b) 72.33
(c) 73.33
(d) 74.33

2.  The elevation change from the PI location on the curve to the PI elevation
is
(a) -27.00
(b) -13.50
(c) +13.50
(d) +27.00

3.  Extend the incoming grade to the PT.  The elevation change from this point
to the actual PT elevation is
(a) -54.00
(b) -27.00
(c) +27.00
(d_ +54.00
4.  The grade at a point 12 stations forward of the PC is
(a) 0%
(b) -1%
(c) -2%
(d) -3%
Answer #1

Rrobu Sony 9,- +2%. ; 9,- -4% ; L= 1800; Sta (PI) = 1800, Ele(P1)= 100 No - Stacp() = Sta(P1)- 4 = 1800-900 = 90o = 9700 Y